# FreqKV: Frequency Domain Key-Value Compression for Efficient Context Window Extension

### v1

@@ 5.2 Long Context Language Modeling
We use FreqKV to train LLaMA-2-base-7b on RedPajama with lengths of 8192, 16384, and 32768. Perplexity is measured on test sets of the book corpus dataset PG-19 and the Arxiv math dataset Proof-pile with the evaluation sliding window of 256. We compare our method with other baselines including full fine-tuning (Full FT), LongLoRA [7], and LoCoCo [5]. While Full FT and LongLoRA leverage full KV cache during inference, LoCoCo, and our FreqKV use compressed cache.
@@
It can be noticed that FreqKV falls behind LongLoRA on Proof-pile at longer context lengths. Proof-pile consists of a diverse set of mathematical texts, including both informal and formal sources, making it particularly challenging for models using compressed caches to predict the next token. In contrast, LongLoRA retains the full KV cache, which naturally benefits next-token prediction in such structured domains. However, the performance of FreqKV does not deteriorate as the context length increases. Our compressed cache method even surpasses Full FT and LongLoRA on PG-19.
@@ 5.3 Long Context Understanding
To further validate the performance on downstream tasks, we use FreqKV to extend the context window of LLaMA-2-chat-7b from 4K to 8K and LLaMA-3-instruct-8b from 8K to 16K. Models are evaluated on the long context understanding benchmark LongBench [3] and Needle-in-a-Haystack [10].
@@
Scores on the 5 categories of LongBench are reported in Table 2. FreqKV consistently improves the performance of LLaMA-2-chat and LLaMA-3-instruct across a majority of these tasks. We compare our method with different KV compression strategies, including LM-Infinite [12], LongHeads [18], SnapKV [17] and PyramidKV [6]. Most baselines use LLaMA-2-chat-7b as their backbone. The retaining ratio is set to 0.5. FreqKV achieves SOTA (state-of-the-art) on 10 long context understanding tasks for LLaMA-2 and 8 tasks for LLaMA-3.
@@
The Needle-in-a-Haystack results are illustrated in Figure 3. FreqKV performs well in extending the context window of LLaMA-2 and LLaMA-3. In contrast, PyramidKV [6] fails beyond the original window boundary. This is because it merely compresses the KV cache at the first decoding step. Out-of-bound position embeddings are used and it prunes tokens after calculating attention weights. Consequently, those unseen positions lead to the performance collapse.
Table 2: Scores of different KV compression methods on LongBench. The superscript “*” means that we reproduce SnapKV and PyramidKV following their official codes for evaluation. The results of LM-Infinite and LongHeads are reported from [18].
Method Single-Document QA Multi-Document QA Summarization Few-shot Learning Code
@@
LLaMA2-chat 18.7 19.2 36.8 25.4 32.8 9.4 27.3 20.8 25.8 61.5 77.8 40.7 52.4 43.8
+LM-Infinite 0.00 18.57 25.33 27.34 31.96 7.76 11.30 2.99 8.72 32.50 29.22 13.82 34.19 24.55
+LongHeads 11.61 22.98 23.76 31.28 24.10 8.87 25.36 20.24 16.18 50.67 79.98 36.74 53.85 44.22
+SnapKV* 17.91 22.9 35.3 26.03 28.21 8.77 24.99 20.94 25.92 64.0 83.36 41.0 60.83 55.14
+PyramidKV* 17.36 23.58 34.89 26.28 28.39 9.17 24.95 20.9 26.04 64.0 83.36 40.92 60.68 55.17
+FreqKV 20.41 21.05 31.2 34.54 34.78 14.47 25.51 21.81 26.9 56.0 84.09 41.53 58.99 58.65
LLaMA3-instruct 22.52 31.83 41.04 44.14 36.68 24.05 28.87 23.25 26.46 75.5 90.23 42.09 56.7 51.84
+PyramidKV* 22.27 31.93 41.03 44.79 36.56 23.86 28.77 23.14 26.62 74.5 90.25 42.53 59.11 53.33
+FreqKV 22.18 42.43 44.19 46.63 40.24 19.57 30.48 23.3 28.53 69.5 88.32 41.73 60.41 51.0
@@
Figure 3: The Needle-in-a-Haystack results, with the x-axis representing the document length (“haystack”) ranging from 1K to 8K tokens for LLaMA-2-chat-7b and 1K to 16K tokens for LLaMA-3-instruct-8b, and the y-axis showing the position of the “needle” (a short sentence) within the document.
To demonstrate the effectiveness of our compression operation, a variant method Dropping is also implemented for comparison as in Figure 3b. Instead of compressing in the frequency domain, Dropping retains the most recent tokens like StreamingLLM [26]. It shares the same training settings as FreqKV. While Dropping suffers from the permanent loss of earlier tokens, FreqKV could retrieve information from a specific sentence (the “needle”) hidden within a document (the “haystack”) and achieve a better performance.
@@ 6.1 Retaining Ratio
We train LLaMA-2-7b on RedPajama [8] with different retaining ratios. The training length is 8K. Models are evaluated on the validation set of PG-19 with the evaluation length ranging from 2K to 64K.
@@ Appendix C Training settings
For long context language modeling, we fine-tune LLaMA-2-base on the RedPajama [8] pre-training dataset for 1000 steps. For long context understanding, we SFT LLaMA-2-chat and LLaMA-3-instruct on the instruction following dataset LongAlpaca [7] for 5 epochs, which consists of 6.28K long-context QA samples.
The total batch size ($GPU\_number\times Batch\_size\_per\_device\times Gradient\_accumulation\_steps$) is 64. The learning rate increases linearly from 1e-6 to 2e-5 with 20 warm-up steps and remains constant in the following steps. The rank used in the LoRA [15] fine-tuning is set to 8. Following LongLoRA [7], the embedding and normalization layers are learnable during training. All the experiments are conducted on ADA6000 and H100 GPUs. Moreover, we equip our method with FlashAttention-2 [9] for further acceleration and memory saving.
@@ Appendix D Compression Overhead
We conduct further studies on the compression overhead of FreqKV. As introduced in Section 4.3, chunk-wise attention is performed for the prefilling tokens. The size of the attention matrix in each chunk is $(N-L-S)\cdot N$ except for the last few tokens. Therefore, the computational cost of self-attention grows approximately linearly with the input length like sliding window attention [4].

### v2

@@ 5.2 Long Context Language Modeling
We use FreqKV to train LLaMA-2-base-7b on RedPajama with lengths of 8192, 16384, and 32768. Perplexity is measured on test sets of the book corpus dataset PG-19 and the Arxiv math dataset Proof-pile with the evaluation sliding window of 256. We compare our method with other baselines including full fine-tuning (Full FT), LongLoRA (Chen et al.,, 2024), and LoCoCo (Cai et al., 024a, ). While Full FT and LongLoRA leverage full KV cache during inference, LoCoCo, and our FreqKV use compressed cache.
@@
It can be noticed that FreqKV falls behind LongLoRA on Proof-pile at longer context lengths. Proof-pile consists of a diverse set of mathematical texts, making it particularly challenging for models using compressed cache to predict the next token. In contrast, LongLoRA retains full KV cache during inference, which naturally benefits next-token prediction in such structured domains. However, the performance of FreqKV does not deteriorate as the context length increases. Our compressed cache method even surpasses Full FT and LongLoRA on PG-19.
@@ 5.3 Long Context Understanding
To further validate the performance on downstream tasks, we use FreqKV to extend the context window of LLaMA-2-chat-7b from 4K to 8K and LLaMA-3-instruct-8b from 8K to 16K. Models are evaluated on the long context understanding benchmark LongBench and Needle-in-a-Haystack.
@@
Scores on the 5 categories of LongBench are reported in Table 2. FreqKV consistently improves the performance of LLaMA-2-chat and LLaMA-3-instruct across a majority of these tasks. We compare our method with different KV compression strategies, including LM-Infinite (Han et al.,, 2024), LongHeads (Lu et al.,, 2024), FastKV (Jo et al.,, 2025), SnapKV (Li et al.,, 2024) and PyramidKV (Cai. et al., 024b, ). Most baselines use LLaMA-2-chat-7b as their backbone. The retaining ratio is set to 0.5. FreqKV achieves SOTA (state-of-the-art) on most of the long context understanding tasks and obtains the highest average score.
@@
The Needle-in-a-Haystack results are illustrated in Figure 3. FreqKV performs well in extending the context window of LLaMA-2 and LLaMA-3. In contrast, PyramidKV fails beyond the original window boundary. This is because it compresses the long prompt after the first decoding step. Out-of-bound position embeddings are used and it prunes tokens after calculating attention weights. Consequently, those unseen positions lead to the performance collapse.
Table 2: Scores of different KV compression methods on LongBench. The superscript “*” means that we reproduce FastKV, SnapKV and PyramidKV following their official codes for evaluation. The results of LM-Infinite and LongHeads are reported from Lu et al., (2024).
Method Single-Document QA Multi-Document QA Summarization Few-shot Learning Code Avg.
@@
LLaMA2-chat 18.7 19.2 36.8 25.4 32.8 9.4 27.3 20.8 25.8 61.5 77.8 40.7 52.4 43.8 35.17
+LM-Infinite 0.00 18.57 25.33 27.34 31.96 7.76 11.30 2.99 8.72 32.50 29.22 13.82 34.19 24.55 19.16
+LongHeads 11.61 22.98 23.76 31.28 24.10 8.87 25.36 20.24 16.18 50.67 79.98 36.74 53.85 44.22 32.13
+FastKV* 18.25 22.46 35.83 28.84 31.34 8.49 24.4 20.96 25.8 60.0 82.67 40.64 57.43 51.52 36.33
+SnapKV* 17.91 22.9 35.3 26.03 28.21 8.77 24.99 20.94 25.92 64.0 83.36 41.0 60.83 55.14 36.81
+PyramidKV* 17.36 23.58 34.89 26.28 28.39 9.17 24.95 20.9 26.04 64.0 83.36 40.92 60.68 55.17 36.84
+FreqKV 20.41 21.05 31.2 34.54 34.78 14.47 25.51 21.81 26.9 56.0 84.09 41.53 58.99 58.65 37.85
LLaMA3-instruct 22.52 31.83 41.04 44.14 36.68 24.05 28.87 23.25 26.46 75.5 90.23 42.09 56.7 51.84 42.51
+PyramidKV* 22.27 31.93 41.03 44.79 36.56 23.86 28.77 23.14 26.62 74.5 90.25 42.53 59.11 53.33 42.76
+FreqKV 23.14 41.6 43.62 46.48 40.54 20.2 29.99 23.67 28.58 69.0 88.95 41.47 60.33 51.11 43.48
@@
Figure 3: The Needle-in-a-Haystack results, with the x-axis representing the document length ranging from 1K to 8K tokens for LLaMA-2-chat-7b and 1K to 16K tokens for LLaMA-3-instruct-8b, and the y-axis showing the position of the “needle” within the document.
To demonstrate the effectiveness of our compression operation, a variant method Dropping is also implemented for comparison as in Figure 3b. Instead of compressing in the frequency domain, Dropping retains the most recent tokens like StreamingLLM (Xiao et al.,, 2024). It shares the same training settings as FreqKV. While Dropping suffers from the permanent loss of earlier tokens, FreqKV could retrieve information from a specific sentence (the “needle”) hidden within a document (the “haystack”) and achieve a better performance.
@@ 6.1 Retaining Ratio
We train LLaMA-2-7b on RedPajama (Computer,, 2023) with different retaining ratios. The training length is 8K. Models are evaluated on the validation set of PG-19 with the evaluation length ranging from 2K to 64K.
@@ Appendix C Training settings
For long context language modeling, we fine-tune LLaMA-2-base on the RedPajama (Computer,, 2023) pre-training dataset for 1000 steps. For long context understanding, we SFT LLaMA-2-chat and LLaMA-3-instruct on the instruction following dataset LongAlpaca (Chen et al.,, 2024) for 5 epochs, which consists of 6.28K long-context QA samples.
The total batch size ($GPU\_number\times Batch\_size\_per\_device\times Gradient\_accumulation\_steps$) is 64. The learning rate increases linearly from 1e-6 to 2e-5 with 20 warm-up steps and remains constant in the following steps. The rank used in the LoRA (Hu et al.,, 2021) fine-tuning is set to 8. Following LongLoRA (Chen et al.,, 2024), the embedding and normalization layers are learnable during training. All the experiments are conducted on ADA6000 and H100 GPUs. Moreover, we equip our method with FlashAttention-2 (Dao,, 2023) for further acceleration and memory saving.
The training time and memory requirements of FreqKV and Full FT on the RedPajama are reported in Table 5. For comparison, we set $Batch\_size\_per\_device$ to 1 and $Gradient\_accumulation\_steps$ to 8. 8$\times$×ADA6000-48GB GPUs are used. Statistics show that FreqKV significantly reduces both training time and memory requirements, demonstrating its efficiency.
Table 5: The training time and memory requirements of FreqKV and Full FT.
Methods Training Length Time (hours) Memory (GB)
Full FT 8K 28.88 44.04
FreqKV 8K 17.07 24.58
16K 38.57 34.00
32K 89.35 45.37
@@ Appendix D Compression Overhead
We conduct further studies on the compression overhead of FreqKV. As introduced in Section 4.3, chunk-wise attention is performed for the prefilling tokens. The size of the attention matrix in each chunk is $(N-L-S)\cdot N$ except for the last few tokens. Therefore, the computational cost of self-attention grows approximately linearly with the input length like sliding window attention (Beltagy et al.,, 2020).